Molecular Physics Vol. 00, No. 00, DD Month 200x, 1–21

Molecular Physics: L^{A}TEX 2ε style guide for authors (Style A)

Taylor & Francis Limited, 4 Park Square, Milton Park, Abingdon, OX14 4RN, UK

(v3.0 released February 2005)

This guide is for authors who are preparing papers for the Taylor & Francis journal Molecular Physics (tMPH ) using the L^{A}TEX 2ε document preparation system and the Class file tMPH2e.cls, which is available via the journal homepage on the Taylor & Francis website (see section 8). Authors planning to submit their papers in L^{A}TEX 2ε are advised to use tMPH2e.cls as early as possible in the creation of their files.

Index to information contained in this guide

1. Introduction 1.1. The tMPH document style 1.2. Submission of L^{A}TEX 2εarticles

to the journal

2. Using tMPH style 2.1. Landscape pages

3. Additional features 3.1. Titles and authors’ names 3.2. Abstracts 3.3. Lists

4. Some guidelines for using standard features 4.1. Sections 4.2. Illustrations (figures) 4.3. Tables 4.4. Running headlines 4.5. Maths environments 4.6. Typesetting mathematics

4.6.1. Displayed mathematics 4.6.2. Bold math italic symbols 4.6.3. Bold Greek

4.6.4. Upright Greek characters 4.7. Appendices 4.8. References

4.8.1. References cited in the text 4.8.2. The list of references

4.9. tMPH macros

5. Example of a section heading with small caps, lowercase, italic, and bold Greek such as κ

6. tMPH journal style 6.1. Punctuation 6.2. Spelling 6.3. Hyphens, n-rules, m-rules and

minus signs 6.4. References 6.5. Maths fonts

7. Troubleshooting 7.1. Fixes for coding problems

8. Obtaining the tMPH2e Class file 8.1 Via the Taylor & Francis website 8.2 Via e-mail

Molecular Physics ISSN 0026-8976 print/ ISSN 1362-3028 online \begin{matrix}c \\ ◯ \end{matrix}2005 Taylor & Francis Ltd http://www.tandf.co.uk/journals DOI: 10.1080/002689700xxxxxxxxxxx

2 L^{A}TEX 2ε guide for authors

1 Introduction

In addition to the standard submission of hardcopy from authors, Molecular Physics (tMPH ) accepts machine-readable forms of papers in Microsoft® Word and L^{A}TEX 2ε. The layout design for tMPH has been implemented as a L^{A}TEX 2ε Class file. The tMPH Class file is based on article.cls. Commands that differ from the standard L^{A}TEX 2ε interface, or which are provided in addition to the standard interface, are explained in this guide. This guide is not a substitute for the L^{A}TEX 2ε manual itself.

1.1 The tMPH document style

The use of L^{A}TEX 2ε document styles allows a simple change of style (or style option) to transform the appearance of your document. The tMPH2e Class file preserves the standard L^{A}TEX 2ε interface such that any document that can be produced using the standard L^{A}TEX 2ε article style can also be produced with the tMPH style. However, the measure (or width of text) is narrower than the default for article, therefore line breaks will change and long equations may need re-formatting.

When your article appears in the print edition of the tMPH journal, it is typeset in Monotype Times. As most authors do not own this font, it is likely that the page make-up will change with the change of font. For this reason, we ask you to ignore details such as slightly long lines, page stretching, or figures falling out of synchronization with their citations in the text, because these details will be dealt with at a later stage.

1.2 Submission of L^{A}TEX 2ε articles to the journal

Papers for consideration should be sent to any of the Editors, as follows: F. Schaefer III (Chair), Center for Computational Quantum Chemistry, University of Georgia, Athens, GA 30602, USA (e-mail: hfs@arches.uga.edu); N. C. Handy, Department of Chemistry, University of Cambridge, Cambridge CB2 1EW, UK (e-mail: nch1@cus.cam.ac.uk); J.-P. Hansen, Department of Chemistry, University of Cambridge, Lensfield Road, Cambridge, CB2 1EW, UK (e-mail: molphys@ch.cam.ac.uk); F. Merkt, Laboratorium für Physikalische Chemie, Eidgenössische Technische Hochschule Zürich ETH-Hönggerberg, HCI CH-8093 Zürich, Switzerland (email: frederic.merkt@mw.phys.chem.ethz.ch); T. P. Softley, Chemistry Research Laboratory, University of Oxford, Oxford, OX1 3TA, UK (e-mail: molecular-physics-editor@chem.ox.ac.uk). Authors are encouraged to submit manuscripts electronically. Electronic submissions should be sent as e-mail attachments using a standard word processing program, such as MSWORD or PDF. If e-mail submission is not possible, please send an electronic version on disc along with three paper copies together with one set of high-quality figures for reproduction. General Instructions for Authors may be found at

(http://www.tandf.co.uk/journals/authors/tmphauth.asp).

Appropriate gaps should be left for figures, of which original versions and copies should also be supplied. Authors should ensure that their figures are suitable (in terms of lettering size, etc.) for the reductions they intend.

Authors who wish to incorporate Encapsulated PostScript artwork directly in their articles can do so by using Tomas Rokicki’s EPSF macros (which are supplied with the DVIPS PostScript driver). See section 2.1, which also demonstrates how to treat landscape pages. Please remember to supply any additional figure macros you use with your article in the preamble before begin{article}. Authors should not attempt to use implementation-specific \special’s directly.

On acceptance (as opposed to initial submission), authors should post the final version of the manuscript to the Editor to whom the initial submission was made, as follows.

(i) a 3.5- or 5.25-inch disc (double- or high-density PC or Apple Mac), containing the corrected version of the paper, plus any macro files you have been using. The final version should be text-only ASCII files (i.e. with no system-dependent control codes);

(ii) two hard copies of the corrected version, plus a marked copy of the accepted version showing any further author corrections incorporated after acceptance ;

(iii) a signed statement that the revised copy and the disc do indeed correspond.

In addition, the disc(s) should be referred to in a covering letter and clearly labelled, as follows.

(i) the disc format (e.g. IBM 360k); (ii) how many files each disc contains, their names, a description of the file

contents and the number of pages each file will produce when printed; details of any user-defined macros;

(iii) the computer system and implementation of L^{A}TEX 2ε used (e.g. IBM AT, PCL^{A}TEX 2ε).

Ensure that any author-defined macros are gathered together in the source file, just before the \begin{document} command.

Please note that, if serious problems are encountered with the coding of a paper (missing author-defined macros, for example), it may prove necessary to divert the paper to conventional typesetting, i.e. it will be re-keyed.

2 Using the tMPH Class file

If the file tMPH2e.cls is not already in the appropriate system directory for L^{A}TEX 2ε files, either arrange for it to be put there, or copy it to your working folder. The tMPH document style is implemented as a complete document style, not a document style option. In order to use the tMPH style, replace

`article' by `tMPH2e' in the \documentclass command at the beginning

of your document:

\documentclass{article}

is replaced by

\documentclass{tMPH2e}

In general, the following standard document style options should not be used with the tMPH style:

(i) 10pt, 11pt, 12pt—unavailable; (ii) oneside (no associated style file)—oneside is the default; (iii) leqno and titlepage—should not be used; (iv) singlecolumn—is not necessary as it is the default style.

2.1 Landscape pages

If a table or illustration is too wide to fit the standard measure, it must be turned, with its caption, through 90^{∘} anticlockwise. Landscape illustrations and/or tables can be produced directly using the tMPH2e style file us-

ing \usepackage{rotating} after \documentclass{tMPH2e}. The following

commands can be used to produce such pages.

\setcounter{figure}{2}
\begin{sidewaysfigure}
\centerline{\epsfbox{fig1.eps}}
\caption{This is an example of figure caption.}
\label{landfig}
\end{sidewaysfigure}
\setcounter{table}{0}
\begin{sidewaystable}
  \tbl{The Largest Optical Telescopes.}
    \begin{tabular}{@{}llllcll}
    .
    .
    .

  \end{tabular}\label{tab1}
\end{sidewaystable}

Before any float environment, use the \setcounter command as above to fix the numbering of the caption. Subsequent captions will then be automatically renumbered accordingly.

3 Additional features

In addition to all the standard L^{A}TEX 2ε design elements, the tMPH style includes separate commands for specifying short versions of the title and authors’ names for running headlines. In general, once you have used this additional tMPH2e.cls feature in your document, do not process it with a standard L^{A}TEX 2ε style file.

3.1 Titles and authors’ names

In the tMPH style, the title of an article and the author’s name (or authors’ names) are used both at the beginning of the article for the main title and throughout the article as running headlines at the top of every page. The title is used on odd-numbered pages (rectos) and the abbreviated list of authors’ names appears on even-numbered pages (versos). Although the main heading can run to several lines of text, the running headline must be a single line (≤ 47 characters). Moreover, the main heading can also incorporate new line commands (e.g. \\) but these are not acceptable in a running headline. To enable you to specify an alternative short title and an abbreviated list of authors’ names, the \markboth command has been used to produce the running headline. The running headlines for this guide were produced using the following code (i.e versos and rectos the same):

\markboth{\LaTeXe\ guide for authors}{\LaTeXe\
guide for authors}

The \thanks note produces a footnote to the title or author(s). In this situation, footnote symbols should be inserted manually in the order: ∗ \ast (for the ‘Corresponding author’), † \dagger,

‡ \ddagger, § \S, ¶ \P, ∥ \|, †† \dagger\dagger,
‡‡ \ddagger\ddagger, §§ \S\S, ¶¶ \P\P, ∥∥ \|\|.

Note that footnotes to the text will automatically be assigned the superscript symbols 1, 2, 3,... by the Class file, beginning afresh on each page.^{1}

1These symbols will be changed to the style of the journal by the typesetter during preparation of your proofs.

The title, author(s) and affiliation(s) should be followed by the \maketitle command.

3.2 Abstracts

Immediately following the \maketitle command, you should include an abstract. The abstract should be enclosed within an abstract environment. For example, the title group and abstract for this guide were produced by the following source code:

\markboth{\LaTeXe\ guide for authors}{\LaTeXe\ guide for authors}
\title{{\textit{Molecular Physics}}:\break
\LaTeXe\ style guide for authors (Style A)}
\author{Taylor \& Francis Limited\thanks{\vspace{6pt}\newline
\centerline{\tiny{{\em Molecular Physics} ISSN 0095-897 print
\textcopyright 2005 Taylor \& Francis Ltd}}
  \newline\centerline{\tiny{
http://www.tandf.co.uk/journals}}\newline \centerline{\tiny{DOI:
10.1080/00958970xxxxxxxxxxxxx}}}, 4 Park Square, Milton Park,
Abingdon, OX14 4RN, UK}
\received{v3.0 released February 2005}
\maketitle
\begin{abstract}
This guide is for authors who are preparing papers for the Taylor \&
Francis journal {\em Molecular Physics} ({\it tMPH}\,) using the \LaTeXe\
document
preparation system and the Class file {\tt tMPH2e.cls}, which is available
via the journal homepage on the Taylor \& Francis website (see section
\ref{FTP}).
Authors planning to submit their papers in \LaTeXe\ are advised to use
{\tt tMPH2e.cls} as early as possible in the creation of their files.
\end{abstract}

3.3 Lists

The tMPH style provides numbered and unnumbered lists using the enumerate environment and bulleted lists using the itemize environment.

The enumerated list numbers each list item with roman numerals:

(i) first item (ii) second item (iii) third item

Alternative numbering styles can be achieved by inserting a redefinition of the number labelling command after the \begin{enumerate}. For example, the list

(1) first item

(2) second item

(3) etc. …

was produced by:

\begin{enumerate}[(2)]
   \item[(1)] first item
   \item[(2)] second item
   \item[(3)] etc. \ldots
\end{enumerate}

Unnumbered lists are also provided using the enumerate environment. For example,

First unnumbered item which has no label and is indented from the left margin. Second unnumbered item. Third unnumbered item.

was produced by:

\begin{enumerate}
  \item[] First unnumbered item...
  \item[] Second unnumbered item.
  \item[] Third unnumbered item.
\end{enumerate}

Itemized lists are provided using the itemize environment. For example,

• First bulleted item

• Second bulleted item

• Third bulleted item

was produced by:

  \begin{itemize}
  \item First bulleted item
  \item Second bulleted item

  \item Third bulleted item
  \end{itemize}

4 Some guidelines for using standard features

The following notes may help you achieve the best effects with the tMPH2e Class file.

4.1 Sections

L^{A}TEX 2ε provides five levels of section headings and they are all defined in the tMPH2e Class file:

(i) \section
(ii) \subsection
(iii) \subsubsection
(iv) \paragraph
(v) \subparagraph

Numbering is automatically generated for section, subsection, subsubsection and paragraph headings. If you need additional text styles in the headings, see the examples in section 5.

4.2 Illustrations (figures)

The tMPH style will cope with most positioning of your illustrations and you should not normally use the optional positional qualifiers of the figure environment, which would override these decisions. See ‘Instructions for Authors’ in the journal’s homepage on the Taylor & Francis website for how to submit artwork. Figure captions should be below the figure itself, therefore the \caption command should appear after the figure. For example, figure 1 with caption is produced using the following commands:

\begin{figure}
\centerline{\epsfbox{fig1.eps}} \caption{This is an example of a
figure caption.}
  \label{sample-figure}
\end{figure}

Figure 1. This is an example of a figure caption.

4.3 Tables

The tMPH style will cope with most positioning of your tables and you should not normally use the optional positional qualifiers of the table environment, which would override these decisions. The table caption appears above the body of the table in tMPH style, therefore the \tbl command should appear before the body of the table.

The tabular environment can be used to produce tables with single thick and thin horizontal rules, which are allowed, if desired. Thick rules should be used at the head and foot only and thin rules elsewhere.

Commands to redefine quantities such as \arraystretch should be omitted. For example, table 1 is produced using the following commands. Note that \rm

Table 1. Radio-band beaming model parameters for FSRQs and BL Lacs.

Class^{a}  γ_{1} γ_{2}^{b} ⟨γ⟩  G  f  θ_{c}
BL Lacs 5  36  7 −4.0 1.0 × 10^{−2} 10^{∘}
FSRQs  5  40 11 −2.3 0.5 × 10^{−2} 14^{∘}

aThis is not as accurate, owing to numerical error.

bAn example table footnote to show the text turning over when a long footnote is inserted.

will produce a roman character in math mode. There are also \bf and \it, which produce bold face and text italic in math mode.

\begin{table}
  \tbl{Radio-band beaming model parameters
           for {FSRQs and BL Lacs.}}
{\begin{tabular}{@{}lcccccc}\toprule
   Class$^{\rm a}$
  & $\gamma _1$ & $\gamma _2$$^{\rm b}$
         & $\langle \gamma \rangle$
         & $G$ & $f$ & $\theta _{c}$ \\
\colrule
   BL Lacs &5 & 36 & 7 & $-4.0$
         & $1.0\times 10^{-2}$ & 10$^\circ$ \\
   FSRQs & 5 & 40 & 11 & $-2.3$
         & $0.5\times 10^{-2}$ & 14$^\circ$ \\
   \botrule
  \end{tabular}}
\tabnote{$^{\rm a}$This is not as accurate, owing to
          numerical error.}
\tabnote{$^{\rm b}$An example table footnote to show the
      text turning over when a long footnote is inserted.}\label{symbols}
\end{table}

To ensure that tables are correctly numbered automatically, the \label{} command should be inserted just before \end{table}.

4.4 Running headlines

As described above, the title of the article and the author’s name (or authors’ names) are used as running headlines at the top of every page. The headline on left-hand pages can list up to two names; for more than two use et al. The \pagestyle and \thispagestyle commands should not be used.

4.5 Maths environments

The tMPH style provides for the following maths environments.

Lemma 4.1 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT.

Theorem 4.2 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT.

Corollary 4.3 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT.

Proposition 4.4 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT.

Proof More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT. □

Remark 1 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT problem.

Algorithm 1 More recent algorithms for solving the semidefinite programming relaxation are particularly efficient, because they explore the structure of the MAX-CUT problem.

These were produced by:

\begin{lemma}
More recent algorithms for solving the semidefinite
programming relaxation are particularly efficient,
because they explore the structure of the MAX-CUT.
\end{lemma}
\begin{theorem}
...
...
\end{theorem}
\begin{corollary}
...

...
\end{corollary}
\begin{proposition}
...
...
\end{proposition}
\begin{proof}
...
...
\end{proof}
\begin{remark}
...
...
\end{remark}
\begin{algorithm}
...
...
\end{algorithm}

4.6 Typesetting mathematics

4.6.1 Displayed mathematics. The tMPH style will set displayed mathematics centred on the measure without equation numbers, provided that you use the L^{A}TEX 2ε standard control sequences open (\[) and close (\]) square brackets as delimiters. The equation

∑_{i=1}^{p} λ_{i} = trace(S) i ∈ ℝ

was typeset in the tMPH style using the commands

\[
  \sum_{i=1}^p \lambda_i = {\rm trace}({\textrm{\bf S}})\qquad
  i\in {\mathbb R}
\].

For those of your equations that you wish to be automatically numbered sequentially throughout the text, use the equation environment, e.g.

∑_{i=1}^{p} λ_{i} = trace(S)  i ∈ ℝ  (1)

was typeset using the commands

\begin{equation}
  \sum_{i=1}^p \lambda_i = {\rm trace}({\textrm{\bf S}})quad
  i\in {\mathbb R}
\end{equation}

4.6.2 Bold math italic symbols. To get bold math italic you can use \bm, which works for all sizes, e.g.

\sffamily
\begin{equation}
   {\rm d}({\bm s_{t_{\bm u}}) = \langle{\bm\alpha({\sf{\textbf L}})}[RM({\bm X}_y
   + {\bm s}_t) - RM({\bm x}_y)]^2 \rangle
\end{equation}
\normalfont

produces

d(s_{t_{u}}) = ⟨α(L)[RM(X_{y} + s_{t}) − RM(x_{y})]^{2}⟩  (2)

Note that subscript, superscript, subscript to subscript, etc. sizes will take care of themselves and are italic, not bold, unless coded individually. \bm produces the same effect as \boldmath. \sffamily...\normalfont allows upright sans serif fonts to be created in math mode by using the control sequence ‘\sf’.

4.6.3 Bold Greek. Bold lowercase as well as uppercase Greek characters can be obtained by {\bm \gamma}, which gives γ, and {\bm \Gamma}, which gives Γ.

4.6.4 Upright lowercase Greek characters and the upright partial deriva-

tive sign. Upright lowercase Greek characters can be obtained with the Class file (which calls up upmath.sty) by inserting the letter ‘u’ in the control code for the character, e.g. \umu and \upi produce μ (used, for example, in the symbol for the unit microns—μm) and π (the ratio of the circumference to the diameter of a circle). Similarly, the control code for the upright partial derivative ∂ is \upartial.

4.7 Appendices

Appendices should be set after the references, beginning with the command \appendices followed by the command \section for each appendix title, e.g.

\appendices
\section{This is the title of the first appendix}
\section{This is the title of the second appendix}

produces

Appendix A: This is the title of the first appendix Appendix B: This is the title of the second appendix

Subsections, equations, theorems, figures, tables, etc. within appendices will then be automatically numbered as appropriate.

4.8 References

4.8.1 References cited in the text. References cited in the text should be quoted by number (e.g. [1], [2,4,10], [21–55], not [21]–[55]). References should be listed in the bibliography in the order in which they first appear within the main text. Each bibliographical entry has a key, which is assigned by the author and used to refer to that entry in the text. In this document, the key ed84 in the citation form \cite{ed84} produces ‘[2]’, and the keys ed84 and aiex02 in the citation form \cite{ed84,aiex02} produce ‘[2, 4]’. The citation for a range of bibliographic entries (e.g. ‘ [2–4, 6]’ )will automatically

be produced by \cite{ed84,lam86,aiex02,aiex00}.

4.8.2 The list of references. The following listing shows some references prepared in the style of the journal:

References

[1] M. Neumann. Mol. Phys., 50, 841 (1983).

[2] D.M.F. Edwards, P.A. Madden, I.R. McDonald. Mol. Phys., 51, 1141 (1984).

[3] R.M. Aiex, M.G.C. Resende, P.M. Pardalos, G. Toraldo. GRASP with path-relinking for the three-index assignment problem. Technical report, AT&T Labs-Research (2000).

[4] R.M. Aiex, M.G.C. Resende, C.C. Ribeiro. Probability distribution of solution time in GRASP: an experimental investigation. Available online at: www.graspintime.com (accessed 23 October [5] F. Glover. In Computing Tools for Modeling, Optimization and Simulation: Interfaces in Com-

puter Science and Operations Research, M. Laguna and J.L. Gonzáles-Velarde (Eds), 2nd edn, pp. 1–24, Kluwer Academic, Boston, MA (2000).

[6] L. Lamport. Efficient algorithms for layer assignment problems. PhD thesis, University of Princeton, NJ (1986).

[7] C.W. Misner (Ed.). Gravitation, pp. 56–79, Freeman, San Francisco, CA(1973).

[8] H. Kern. The resurgent Japanese economy and a Japan–United States free trade agreement, paper presented at the 4th International Conference on the Restructuring of the Economic and Political System in Japan and Europe, Milan, Italy, 21–25 May (1996).

[9] F. French. Title of a Book in Another Language, P. Smith, Translator, Dover, New York (1988, original work published 1923).

This list was produced by:

\begin{thebibliography}{9}
\bibitem{neu83}%1
M. Neumann. {\itshape Mol. Phys.,} {\bfseries 50}, 841 (1983).
\bibitem{ed84}%2
D.M.F. Edwards, P.A. Madden, I.R. McDonald. {\itshape Mol. Phys.,}
{\bfseries 51}, 1141 (1984).
\bibitem{aiex00}%3
R.M. Aiex, M.G.C. Resende,  P.M. Pardalos, G. Toraldo. GRASP with
path-relinking for the three-index assignment problem. Technical
report, AT\&T Labs-Research (2000).
\bibitem{aiex02}%4
R.M. Aiex, M.G.C. Resende,  C.C. Ribeiro. Probability distribution
of solution time in GRASP: an experimental investigation.
Available online at: www.graspintime.com (accessed 23 October
2003).
\bibitem{glov00}%5
F. Glover. In {\it Computing Tools for Modeling, Optimization and
Simulation: Interfaces in Computer Science and Operations
Research}, M. Laguna and J.L. Gonz\'{a}les-Velarde (Eds), 2nd edn,
pp.~1--24, Kluwer Academic, Boston, MA (2000).

\bibitem{lam86}%6
L. Lamport. Efficient algorithms for layer assignment problems.
PhD thesis, University of Princeton, NJ (1986).
\bibitem{mtw73}%7

Table 2. tMPH macros

\markboth{optional short author(s)}  short author(s) list and short title used
{optional short title}  in running heads (verso/recto, resp.)
\thanks{title-page footnote to article title  e.g. ‘Corresponding author. E-mail:
or author}  A.N. Author@uiowa.edu’
\begin{abstract}...\end{abstract}  for abstract on titlepage
\bm{math and symbols}  bold italic math and symbols
\bi{text}  bold italic text
\sf{text or upright symbols in math mode} sans serif text or upright symbols in math mode

C.W. Misner (Ed.). {\itshape Gravitation}, pp. 56--79, Freeman,
San \nobreak Francisco, CA(1973).
\bibitem{hk96}%8
H. Kern. The resurgent Japanese economy and a Japan--United States
free trade agreement, paper presented at the 4th International
Conference on the Restructuring of the Economic and Political
System in Japan and Europe, Milan, Italy, 21--25 May (1996).
\bibitem{fzf88}%9
F. French. {\itshape Title of a Book in Another Language}, P.
Smith, Translator, Dover, New York (1988, original work published
1923).
\end{thebibliography}

Each entry takes the form:

\bibitem{key} Bibliography entry

where key is the tag that is to be used as an argument for the \cite{} command. Bibliography entry should be the material that is to appear in the bibliography, suitably formatted.

4.9 tMPH macros

Table 2 gives a list of macros for use with tMPH. The list displays each macro’s code and a description/demonstration of its function.

5 Example of a section heading with small caps, lowercase, italic, and bold Greek such as κ

The following code shows how to achieve this section head:

\section{Example of section heading with\\*
   {\fontencoding{T1}\scshape\lowercase{small caps}},
   \lowercase{lowercase}, {\bi italic},
   and bold\\* Greek such as
   ${\bm\kappa}$}\label{headings}

6 tMPH journal style

The notes given here relate to common style errors found in tMPH manuscripts, but are not intended to be exhaustive.

6.1 Punctuation

When deciding where to add commas, it may be helpful to read through the sentence and note where the natural ‘pauses’ occur. The needs of readers for whom English is not a first language should be borne in mind when punctuating long sentences. For example, consider the following sentence as it appeared in tMPH: ‘When we do not limit ourselves by constraints arising from the choice of an initial fluctuation spectrum, structures in an open universe, including the peculiar velocity structure, can be reproduced in a flat Lemaître universe for a large part of their evolution.’ Now consider the same sentence without commas: ‘When we do not limit ourselves by constraints arising from the choice of an initial fluctuation spectrum structures in an open universe including the peculiar velocity structure can be reproduced in a flat Lemaître universe for a large part of their evolution.’

6.2 Spelling

Please use British spelling—e.g. centre not center, labelled not labeled. The following style regarding -ise, -yse and -ize spellings is used: -ise—devise, surprise, comprise, revise, exercise; -yse—analyse; -ize: recognize, criticize, minimize, emphasize, organize.

6.3 Hyphens, n-rules, m-rules and minus signs

(i) Hyphens (one dash in TEX/L^{A}TEX 2ε). tMPH uses hyphens for compound adjectives (e.g. low-density gas, least-squares fit, two-component model) but not for complex units or ranges, which could become cumbersome (e.g. 15 km s^{−1} feature, 100–200 μm observations).

(ii) n-rules (two dashes in TEX/L^{A}TEX 2ε). These are used (a) to denote a range (e.g. 1.6–2.2 μm); and (b) to denote the joining of two words of equal standing (e.g. Kolmogorov–Smirnov test, Herbig–Haro object).

(iii) The m-rule (three dashes in TEX/L^{A}TEX 2ε) is used in tMPH as an alternative to parentheses (e.g. ‘the results—assuming no temperature gradient—are indicative of … ’).

(iv) The minus sign (one dash in TEX/L^{A}TEX 2ε) is produced automatically in math mode by use of a single dash, e.g.

y_{i} ∈ {−1, 1} ∀i ∈ V  (3)

where | − V | = A^{2} + B^{2}

is produced by

\begin{equation}
y_{i} \in \{-1, 1 \} \quad \forall i \in V
\end{equation}
\noindent where $|-V|=A^2+B^2$

6.4 References

It is important to use the correct reference style, details of which can be found in section 4.8 above.

6.5 Maths fonts

Scalar variables should be mediumface italic (e.g. s for speed); vectors should be bold italic (e.g. v for velocity); matrices should be bold roman (upright) (e.g. A), and tensors should be bold upright sans serif (e.g. L). Differential d, partial differential ∂, complex i, exponential e, superscript T for ‘transpose’, sin, cos, tan, log, etc., should all be roman. Openface, or ‘blackboard’, fonts can be used, for example, for the integers ℤ and the reals ℝ. Sub/superscripts that are physical variables should be italic, while those that are labels should be roman (e.g. C_{p}, T_{eff}). Displayed equations should have end-of-line punctuation appropriate to the running text sentence of which they form a part.

7 Troubleshooting

Authors may from time to time encounter problems with the preparation of their papers in L^{A}TEX 2ε. The appropriate action to take will depend on the nature of the problem – the following is intended to act as a guide.

(i) If the problem is with L^{A}TEX 2ε itself, rather than with the actual macros, please refer to the appropriate handbooks for initial advice.^{1} If the solution cannot be found, and you suspect that the problem lies with the macros, then please contact Taylor & Francis (latex.helpdesk@tandf.co.uk).

(ii) Problems with page make-up (e.g. large spaces between paragraphs, or under headings or figures; uneven columns; figures/tables appearing out of order): please do not attempt to remedy these yourself using ‘hard’ page make-up commands – the typesetter will correct such problems. (You may, if you wish, draw attention to particular problems when submitting the final version of your paper.)

(iii) If a required font is not available at your site, allow TEX to substitute the font and specify which font your require in the covering letter accompanying your file(s).

7.1 Fixes for coding problems

This guide has been designed to minimize the need for user-defined macros to create special symbols. Authors are urged, wherever possible, to use the following coding rather than to create their own. This will minimize the danger of author-defined macros being accidentally ‘over-ridden’ when the paper is typeset in Times (see section 4.6, ‘Typesetting mathematics’ above). In cases where it is essential to create your own macros, these should be displayed in the preamble of the source file before \begin{document}.

(i) Fonts in section headings and paper titles. The following are examples of styles that sometimes prove difficult to code.

Paper titles:

Generalized Flory theory at δ > 50^{∘}

is produced by

\title{Generalized Flory theory at

1TEX: Knuth, D., 1986, The TEX book (New York: Addison–Wesley); L^{A}TEX 2ε: Lamport, L., 1985, L^{A}TEX 2ε User’s Guide and Reference Manual (New York: Addison–Wesley).

     ${\bm\delta > {\bfseries 50}^\circ}$}

Ion–ion correlations in H ii regions

is produced by

\title{Ion--ion correlations in H\,{\sc ii} regions}

(ii) n-rules, m-rules, hyphens and minus signs (see section 6.3 for correct usage). To create the correct symbols in the sentence

The high-resolution observations were made along a line at an angle of −15^{∘} (East from North) from the axis of the jet—which runs North–South

you would use the following code:

The high-resolution observations were made along a line at an
angle of $-15^\circ$ (East from North) from the axis of the
jet---which runs North--South

(iii) Fonts in superscripts and subscripts. Subscripts and superscripts will automatically come out in the correct font and size in a math environment (e.g. enclosed by ‘$’ delimiters in running text or within \[...\] or the ‘equation’ environment for displayed equations). You can create the output k_{x} by typing ${\bm k_x}$. If the subscripts or superscripts need to be other than italic, they should be coded individually—see (vi) below.

(iv) Calligraphic letters (uppercase only). Normal calligraphic can be produced

with \cal as usual (in math mode).

(v) Automatic scaling of brackets. The codes \left and \right should be used to scale brackets automatically to fit the equation being set. For example, to get

v = x (\frac{N + 2}{N})

use the code

\[
   v = x \left( \frac{N+2}{N} \right)
\]

(vi) Roman font in equations. It is often necessary to make some symbols roman in an equation (e.g. units, non-variable subscripts). For example, to get the following output:

σ ≃ (r/13 h^{−1} Mpc)^{−0.9}, ω = \frac{N − N_{s}}{N_{R}}

you should use:

   \[
      \sigma \simeq (r/13~h^{-1}
      ~{\rm Mpc})^{-0.9}, \qquad \omega
      =\frac{N-N_{{\rm s}}}{N_{{\rm R}}}
   \]

8 Obtaining the tMPH2e Class file

8.1 Via the Taylor & Francis website

This Guide for Authors and the tMPH2e.cls Class file may be obtained via the Instructions for Authors on the Taylor & Francis homepage for the journal

(http://www.tandf.co.uk/journals/titles/00268976.asp).

Please note that the Class file calls up the following open-source LaTeX packages, which will, for convenience, unpack with the downloaded Guide for Authors and Class file: amsfonts.sty; amssymb.sty; cite.sty; rotating.sty; and upmath.sty.

8.2 Via e-mail

This Guide for Authors, the Class file and the associated open-source La-TeX packages are also available by e-mail. Requests should be addressed to latex.helpdesk@tandf.co.uk clearly stating for which journal you require the Guide for Authors and/or Class file.